Session: Nov 2004 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Thu Oct 28 09:24:32 2004) Success :
The label number of the mark5 discs is SHAO-012. Please produce the last four codes of the label, because of we haven't the code scanner. And I can't use the track for the discs. It will be send out from Shanghai tomorrow.

Hartebeesthoek (Jonathan Quick at Thu Oct 28 11:14:56 2004) Success :
No problems

Noto (P.Cassaro at Thu Oct 28 11:23:05 2004) Success :

Urumqi (Liu Xiang at Thu Oct 28 11:26:40 2004) Minor failures : WIND HIGH
Snow and rain just before the experiment, and then it is wet and foggy, so the Tsys is very high about 160K that time, and then it is slowly decreased to about 30K at the end.

Onsala (Michael Lindqvist at Thu Oct 28 11:36:09 2004) Success :
No known problems.

Torun (K. Borkowski at Thu Oct 28 11:55:12 2004) Success :
No known problems.

Medicina (Cattani A. at Thu Oct 28 12:22:59 2004) Success :

Effelsberg (D.Graham at Fri Oct 29 08:25:18 2004) Minor failures :
looking into hills at start, high tant of 3c273 also causes problems. Otherwise OK

Jodrell_Bank (Alastair Gunn at Mon Nov 1 00:59:24 2004) Success :
The Lovell Telescope was used. No known problems.

Westerbork (Tony Foley at Mon Nov 1 11:10:29 2004) Success :
Wb used in phased array mode. No known problems

Urumqi (Liu Xiang at Tue Nov 2 08:34:17 2004) Success :
We don't want to send disk-pack for n04c3 because we lack of disk-pack and we got ftp fringes already.